Solve:
(4x - 5y = 10
(3x - 2y = 4 (Use elimination.)

Answer by checkley77(12569)

4x - 5y = 10 multiply by -3
3x - 2y = 4 multiply by 4 & then add.
-12x+15y=-30
12x-8y=16
------------------------------------------------------
7y=-14
y=-14/7
y=-2 answer.
4x-5*-2=10
4x+10=10
4x=10-10
4x=0
x=0 answer.